### v2

# Physics Priors Offer Useful Accuracy-Carbon Trade-Offs in Spatio-Temporal Forecasting

Abstract
Development of modern deep learning methods has been driven primarily by the push for improving model efficacy (accuracy metrics). This sole focus on efficacy has steered development of large-scale models that require massive computational resources, and results in considerable energy consumption and corresponding carbon footprint across the model lifecycle. In this work, we explore how physics inductive biases can offer useful trade-offs between model efficacy and model efficiency (compute, energy, and carbon). We study models with strong, weak, and no physics-inductive biases for spatio-temporal forecasting of incompressible shear flow, a task governed by the Navier-Stokes equations. We find that models with stronger physics priors achieve substantially lower training footprints, but this advantage does not straightforwardly extend to inference, highlighting the importance of evaluating carbon costs across the full model lifecycle rather than any single stage. We argue that model efficiency, along with model efficacy, should become a core consideration driving machine learning model development and deployment.
1 Introduction
Machine learning (ML) has advanced rapidly over the past decade, driven largely by the pursuit of improving model efficacy (accuracy metrics). This narrow focus has led to increasingly large models that require substantial computational resources and generate significant carbon emissions across their lifecycle [1, 2, 3, 4]. As models scale, efficacy gains diminish while costs escalate, underscoring the need to move beyond accuracy as the sole measure. A more holistic assessment must balance efficacy with efficiency, capturing not only predictive performance but also compute, energy, and carbon costs [5]. This requires identifying strategies that strike a favourable balance rather than optimising for accuracy alone.
Figure 1: a: Conceptual map of modelling approaches by data availability (vertical axis) and domain knowledge (horizontal axis). Physics-informed ML lies in the middle, combining data-driven flexibility with physics priors. b: Test mean squared error (MSE) versus CO2eq emissions for a data-driven NN (red squares), an unsupervised PINN (blue dots), and a semi-supervised PINN (grey triangles). The dashed line marks the Pareto front, dominated by the S-PINN, indicating the best accuracy-carbon trade-off. Details of the experiment are in App. 0.A.
Physics-informed ML offers a promising path towards this balance. By embedding prior knowledge about the underlying system directly into the model, one can combine the flexibility of data-driven methods with the structure of domain knowledge, as illustrated in Fig. 1-a. Prior work has shown that embedding conservation laws, symmetries, or governing equations can reduce the number of trainable parameters [6, 7], reduce the amount of training data [8, 9, 10], and accelerate training convergence [11, 12].
Whether physics-informed ML models can also reduce energy consumption and carbon footprint remains a less explored area. Fig. 1-b offers a concrete illustration of its potential. In a simple experiment on a partial differential equation (PDE) system, we compare three models that share the same underlying architecture but differ in how much they rely on physics through their loss functions: a purely data-driven neural network (NN), an unsupervised physics-informed neural network (PINN) trained solely on PDE residuals, and a semi-supervised PINN combining data and physics. The semi-supervised PINN dominates the Pareto front, achieving lower prediction error at a lower carbon cost than either alternative. This suggests that striking the right balance between data and physics does not just improve accuracy, but can simultaneously reduce the carbon footprint. Full details of this experiment are provided in App. 0.A.
Building on this intuition, we extend the perspective from task-specific models such as PINNs to general-purpose architectures that embed physics inductive biases in ways that are transferable across datasets and domains. We focus on spatio-temporal forecasting of PDE dynamics, a setting characterised by high computational cost and practical relevance. Operational weather and climate forecasting, for instance, requires large-scale PDE solvers running on supercomputers multiple times per day [13]. In such contexts, even modest efficiency improvements can translate into substantial operational carbon savings [14], making this domain an ideal test-bed for exploring sustainable ML design [15]. Our key contributions towards this end are:
  1. 1.
    Carbon-aware evaluation: We promote evaluation practices that explicitly account for training and inference costs alongside predictive accuracy.
  2. 2.
    Characterising the accuracy-carbon trade-off: We provide an empirical characterisation of how varying degrees of physics priors affect the accuracy-carbon trade-off in spatio-temporal forecasting, evaluated on a complex, high-resolution dataset of incompressible shear flow.
2 Background and Related Work
Figure 2: Bias-variance trade-off. Top: classical trade-off with minimum total error at optimal model complexity. Bottom: possible analogous balance between physics-based constraints and data-driven flexibility.
Revisiting the Bias-Variance Trade-Off.
It is commonly understood in ML that strong model assumptions (bias) risk under-fitting, whereas high sensitivity to training data can lead to over-fitting (variance) [16]. Striking the right balance is crucial to obtain the right class of models.
We draw an analogy to model design under resource constraints. Incorporating domain knowledge as inductive bias can reduce reliance on large datasets and potentially lower computational demands, but may limit model flexibility. Conversely, highly data-driven models offer flexibility, but must learn all patterns from data which often requires lots of data and computational resources. We therefore hypothesise that a physics-data balance may exist, analogous to the bias-variance trade-off (Fig. 2).
The Question of Resource-Awareness.
The standard metrics used to characterise the performance of ML models primarily focus on efficacy and are untethered from efficiency considerations. This introduces a preference for models that could offer high efficacy at large resource costs [17]. There are two primary streams of work that aim to include resource-aware metrics: using composite metrics and multi-objective optimisation.
Several composite metrics have been proposed to combine predictive performance with resources like energy consumption or carbon footprint. Evchenko et al. [18] introduced a composite metric that combines efficacy and efficiency metrics to study the influence of resource constraints on traditional (non deep learning) ML models. Carburacy [19] extends the idea to transformer models by jointly quantifying accuracy and carbon emissions. More recently, Kapoor et al. [20] proposed EcoL2, a composite metric similar to Carburacy but extended to include the costs of data generation and hyperparameter tuning.
Multi-objective optimisation has been primarily used in neural architecture search (NAS) resulting in models like EfficientNet [21]. More recently, energy consumption [17] and carbon footprint [22] have been taken into consideration in NAS. These methods, however, have not actively studied the influence of physics inductive biases on the resource consumption of ML models.
Spatio-Temporal Forecasting Models.
There are many model classes for the task of spatio-temporal forecasting, differing in the extent of physics inductive bias they use. In this work, we focus on two: the U-net [23] and the Fourier Neural Operator (FNO) [24].
U-nets perform forecasting by mapping a window of past states to future states. They capture local correlations by fusing features across multiple spatial scales, but do not impose any explicit constraints on the underlying dynamics beyond those learning from data. Mild forms of inductive biases can be introduced to convolutional architectures like U-net, for example through circular padding to respect periodic boundaries and avoid edge discontinuities.
The FNO incorporates a stronger inductive bias through its Fourier-space representation, enabling the model to capture global spectral interactions. This makes it particularly well-suited to systems with periodicity and long-range correlations. In our setting, we adapt the FNO from its original operator-learning formulation [25] to an autoregressive setup, where its spectral parameterisation explicitly encodes global and periodic structure, in contrast to the primarily local bias of the U-net.
3 Experiment
Problem Setup.
We consider spatio-temporal forecasting of dynamical systems, where the objective is to approximate the solution $u(x,y,t)$ of a PDE given initial and boundary conditions. We frame this as an autoregressive prediction task on a two-dimensional, uniformly spaced mesh $(x,y)\in\{x_{1},x_{2},\ldots,x_{H}\}\times\{y_{1},y_{2},\ldots,y_{W}\}$ with discretized time steps $t\in\{t_{1},t_{2},\ldots,t_{T}\}$. At each step $t_{k}$ the state is represented as a tensor $u_{k}(x,y)\in\mathrm{R}^{H\times W\times C}$, where $k\in\{1,2,..,T\}$ indexes discrete time steps in a trajectory of length $T$. $H$ and $W$ are spatial dimensions and $C$ is the number of physical fields (channels), such as pressure or velocity components.
To capture temporal dependencies, the model is conditioned on a finite history of length $h$. The input window is denoted as $u_{k-h+1:k}(x,y):=(u_{k-h+1}(x,y),\\ \ldots,u_{k}(x.y))$. A parameterised function $f_{\theta}$ is trained to predict:
$u_{k+1}(x,y)=f_{\theta}(u_{k-h+1:k}(x,y)\big).$ (1)
During training, the loss is computed on one-step predictions. At inference, we assess long-term stability via autoregressive rollouts, obtained by iteratively applying Eq. 1 and feeding the model’s own predictions back as inputs.
We instantiate this framework for forecasting incompressible shear flow governed by the Navier-Stokes equations, a setting that is both computationally demanding and governed by well-understood physical structure, making it well-suited for studying the efficacy-efficiency trade-off.
Data.
We use the incompressible shear flow data from The Well collection [26], generated with Dedalus [27]. The dataset captures the coupled non-linear evolution of four physical fields: horizontal and vertical velocity, pressure, and a passive tracer on a periodic spatial domain, see example in Fig. 3. We work with a subset, which consists of 240 simulations spanning six PDE parameter settings and $40$ initial conditions settings. Simulations are resolved at $256\times 512$ with $200$ time steps ($\Delta t=0.1$). The data is split by initial condition into 8080/1010/$10$ for training, validation, and testing, yielding $37,632$ one-step samples for training, $4,704$ for one-step inference, and $24$ trajectories for autoregressive rollout initialised at $t=0$. Further details on the PDE formulation and initialisation are provided in App. 0.B.
Figure 3: Example of incompressible shear flow data. Temporal evolution of tracer ($s$), pressure ($p$), horizontal velocity ($v_{x}$) and vertical velocity ($v_{y}$) fields at four time steps: $t\in\{0,T/3,2T/3,T\}$. The snapshots illustrate the coupled dynamics of the fields.
Table 1: Overview of the seven models and the relative strength of their physics inductive biases.
Models Physics inductive bias Strength
UNet, UNet-CN None None
UNet-CP, UNet-CN-CP Circular padding Weak
FNO, TFNO, UFNO Spectral layers Strong
Models.
We evaluate seven models spanning a spectrum of physics inductive biases, summarised in Table 1. We consider a baseline U-net and a ConvNeXt variant (UNet-CN) [28], which improves hardware efficiency through modern convolutional design. We introduce a weak form of inductive bias by applying circular padding to both architectures (UNet-CP, UNet-CN-CP). For models with stronger physics priors, we include three FNO variants: a baseline FNO, a Tucker-factorized variant (TFNO) [29] that improves parameter efficiency, and a U-shaped variant (UFNO) [30] designed to better capture multi-scale interactions.
Training Procedure.
All models use a history of $h=4$ states and are trained for one-step prediction according to Eq. 1. Across all models, we use a batch size of 16 and the AdamW optimizer [31] with $10^{-4}$ weight decay. We adopt the coarse-tuned learning rates reported in The Well [26] ($5\cdot 10^{-4}$ for U-nets and $10^{-3}$ for FNOs, respectively), which were tuned across several datasets including the one used here. A linear warm-up cosine scheduler was applied for the first three epochs and the loss is the mean squared error (MSE) averaged across fields and space. Training ran on a single Nvidia A40 GPU ($\leq$ 24h) with early stopping (patience 6). Full model configurations are listed in App. 0.C.
Carbon Footprint.
To estimate energy consumption during training and inference, we employ Carbontracker [2]. We report results in terms of carbon footprint rather than energy consumption directly, since energy is only problematic to the extent that it is generated from carbon-emitting sources. Were the grid entirely renewable, the same energy expenditure would carry negligible direct carbon emissions. In practice, however, the global energy mix remains far from carbon-neutral, and we therefore convert energy to carbon dioxide equivalent (CO2eq) emissions using the global average carbon intensity for 2024 at $445$ gCO2/kWh [32]. Using a fixed carbon intensity ensures fair comparisons across models, independent of when or where they were trained.
4 Results
Figure 4: Pearson r over 20 rollout steps. Values are averaged over trajectories and fields. Models labels are ordered from no bias (top) to strong bias (bottom).
Predictive Performance.
[…]
Carbon Footprint.
We next evaluate the carbon footprint and runtime associated with both training and inference (Table 2). Considering carbon efficiency (CO2eq per epoch), FNO and TFNO are roughly as carbon-intensive as the baseline UNet, while the other U-net variants are considerably more intensive. Within the U-net family, both the ConvNext design and circular padding make the model more carbon-intensive, with the exception of UNet-CN-CP, which is more efficient relative to UNet-CN. This indicates that architectural design choices can interact with computational efficiency in non-trivial ways. UFNO stands out as the by far most carbon-efficient model.
Table 2: Carbon footprint during training and inference. Reported CO2eq emissions and runtime for all models, measured with CarbonTracker. Inference results correspond to one-step forecasts over the full test set (4,704 samples) and include the initial model loading cost. Each inference experiment was repeated three times; CO2eq is reported as mean and standard deviation, and runtimes as mean values.
Training Inference (one-step)
Model CO2eq (kg) Time (h) # Epochs CO2eq/Epoch (g) CO2eq (g) Time (min)
UNet $2.72$ $12.7$ $43$ $63$ $3.68\pm 0.03$ $1.0$
UNet-CN $4.71$ $15.9$ $27$ $174$ $7.77\pm 0.05$ $1.9$
UNet-CP $3.81$ $14.2$ $35$ $109$ $4.06\pm 0.05$ $1.0$
UNet-CN-CP $5.89$ $20.9$ $45$ $131$ $4.87\pm 0.02$ $1.2$
FNO $0.80$ $4.0$ $12$ $67$ $7.87\pm 0.94$ $2.0$
TFNO $1.09$ $5.3$ $16$ $68$ $7.21\pm 0.12$ $1.8$
UFNO $1.12$ $3.6$ $30$ $37$ $9.00\pm 0.88$ $2.4$
Figure 5: a: Training emissions (kgCO2eq) from a single run per model. b: One-step inference emissions (gCO2eq) from three repeated runs over the full test set; error bars indicate variability across repetitions. Note that inference costs also include the initial model loading cost. c: Carbon footprint of inference during rollouts. Reported CO2eq emissions for 24 autoregressive rollouts of 50 steps. Inference costs include the initial model loading cost. Models are ordered from no bias (left) to strong bias (right).
Fig. 5 summarises emissions for training (a), one-step inference (b), and rollout inference (c). Comparing total training emissions, the carbon footprints of the FNO variants are between 2.4 and 7.4 times lower than those of the U-net variants. For one-step inference the pattern changes. Here, the FNO variants are on average more expensive than the U-nets. During rollout inference, however, the differences between models become smaller. This is likely because the fixed overhead of model loading constitutes a larger fraction of the total cost. These discrepancies highlight the importance of considering both training and inference costs rather than focusing on a single stage of the modelling pipeline. Although carbon footprint and runtime are naturally correlated, they are not strictly proportional across models (Fig. 6).
5 Discussions
Figure 6: Training runtime versus carbon footprint. The plot shows that energy consumption does not scale one-to-one with runtime.
Fig. 6 emphasises the importance of explicit energy or carbon measurement. Runtime alone is an unreliable proxy, as different hardware operations have distinct power profiles and a model that trains faster does not necessarily train cheaper [5].
Our study demonstrates that physics inductive biases impact the trade-off between efficacy and efficiency. In the simple PINN experiment (Sec. 1), combining data with physics yielded the model dominating the Pareto front, achieving lower test errors at lower carbon cost compared to both the purely data-driven and purely physics-driven baselines.
In the incompressible shear flow setting, we observe clear differences in how the physics priors affect both predictive performance and carbon footprint across training and inference. Incorporating weak physics-inductive biases in the form of periodic padding, improves performance but increases total training emission, while its effects on training carbon efficiency and one-step inference are model-dependent: increasing both in one case (UNet-CP) and decreasing both in the other (UNet-CN-CP). Models with stronger physics inductive biases achieve the lowest training footprints, but this comes at the expense of higher inference costs compared to U-net variants. UFNO achieves one of the most favourable overall profiles, combining strong predictive accuracy with a low training footprint. Its inference cost, however, remains nearly twice that of the best-performing U-net variant (UNet-CN-CP). The results underscore the importance of evaluating training and inference costs together. A model that trains cheaply may prove inefficient when repeatedly deployed, while others entail higher upfront cost but amortise more effectively in deployment. Amortising development costs across the model lifecycle is one way to factor in these considerations [33].
Several key points follow from these results. First, model selection cannot be based on efficacy alone, as models with similar predictive accuracy can differ substantially in carbon emissions. Second, once both training and inference costs are accounted for, identifying a single “best” model becomes less straightforward. A model with a small carbon footprint during training may be unsuitable in deployment, making carbon-aware evaluation across the full lifecycle essential.
The implications extend well beyond our test-bed. Many scientific and engineering problems are governed by well-understood structures, and overlooking them is wasteful. At the same time, tailoring models too closely to the specific physics can limit applicability and can be challenging for complex systems (e.g., PINNs). This motivates the development of general-purpose models such as FNOs, which allow physics priors to be incorporated more easily while retaining flexibility across tasks. However, alignment with the underlying physics does not guarantee carbon efficiency. Physics should not be incorporated at any cost, but only where it provides clear benefits in terms of predictive performance or carbon efficiency. While our work centres on physics-inductive biases, the same logic extends to other forms of prior knowledge: any regularising bias that constrains the learning task appropriately can potentially offer similar trade-offs between efficacy and efficiency.
Ultimately, our results argue for a shift in mindset. Progress in ML should not be equated with ever-larger datasets, models, and compute budgets, but rather evaluated in terms of the value gained relative to the cost incurred [34]. Learning everything from data while ignoring domain knowledge reflects an abundance mindset, in which resources are treated as unlimited. A more sustainable alternative is to exploit what we already know by embedding this knowledge into models. This work offers a step towards understanding how prior knowledge shapes the efficacy-efficiency trade-offs, and argues for evaluating these explicitly.
Limitations.
We did not perform full hyperparameter sweeps for the models. As detailed in App. 0.C, all configurations follow settings from prior work tuned on several PDE datasets, including the shear-flow dataset used here. This may influence the observed trade-offs. Furthermore, we restricted our study to spatio-temporal forecasting. While this limits the scope of the contribution, the central message about taking efficiency considerations into account by including appropriate inductive biases is applicable across ML tasks. Finally, the evaluation metrics focus on predictive accuracy rather than physical fidelity. Physics-based metrics would more directly assess whether the models capture underlying structures or conservation properties.
6 Conclusions
The primary message in this work, backed by empirical evidence, is to increase resource-awareness when designing ML experiments. While this can be accomplished in several ways [35], we focus on the role of physics inductive biases and illustrate the different decisions practitioners have to consider. Using spatio- temporal forecasting of incompressible shear flow, we show how physics-inductive biases affect the trade-off between predictive performance and carbon footprint.
This work advances a perspective grounded in resource-awareness and calls for a shift in mindset. As a community, we should discourage the pursuit of marginal improvements in performance that come from disproportionately higher resource costs. With our work we have made the case to encourage resource-awareness by utilising existing domain knowledge and/or other inductive biases.
{credits}
6.0.1 Acknowledgements
SW and RS acknowledge funding received from Independent Research Fund Denmark (DFF) under grant agreement number 4307-00143B. JHC acknowledges funding by DFF through the project Arctic Push, grant number 4258-00025B. RS also acknowledges funding received under European Union’s Horizon Europe Research and Innovation Action programme under grant agreements No. 101070284, No. 101070408 and No. 101189771.
The authors thank Erik B Dam, Pedram Bakhtiarifard, Yijun Bian and others from the Machine Learning Section at UCPH for constructive feedback on the manuscript. The authors thank members of SAINTS Lab for useful discussions throughout.
[…]
The authors have no competing interests to declare that are relevant to the content of this article.
Appendix
Appendix 0.A Simple Experiment with PINNs
Figure 7: Target and training data for the viscous Burgers’ equation. a: Target data. b: Supervised samples. c: PDE residual samples. d: Initial and boundary condition points.
0.A.1 Experimental Setup
In this simple experiment, we explore PINNs [36] which embed governing PDEs directly into the training objective, thereby regularising the optimisation landscape and guiding solutions towards physical consistency. By minimizing PDE residuals, PINNs can be trained with little or no supervised data.
We study the solution of the viscous Burgers’ equation with Dirichlet boundary conditions on a 2D spatio-temporal grid:
$\frac{\partial u}{\partial t}+u\frac{\partial u}{\partial x}-\nu\frac{\partial^{2}u}{\partial x^{2}}=0,$ (2)
$u(x,0)=-\sin(\pi x),$ (3)
$u(1,t)=u(-1,t)=0,$ (4)
for $x\in[-1,1]$, $t\in[0,1]$, and with fixed viscosity ($\nu=0.025$). The domain is discretized with $256$ spatial points and $100$ time steps (Fig. 7).
We compare three models that share the same underlying architecture (four hidden layers, width 50, tanh activations) but differ in their use of physics through the loss function. The purely data-driven NN minimizes the mean-squared-error (MSE) between predictions $\hat{u}$ and ground truth $u$:
$\mathcal{L}_{\text{NN}}=\mathcal{L}_{\text{data}}=\frac{1}{N_{\text{data}}}\sum_{i=1}^{N_{\text{data}}}(u_{i}-\hat{u}_{i})^{2}.$ (5)
The unsupervised U-PINN is trained solely on physics constraints, minimizing PDE residuals alongside deviations from initial and boundary conditions:
$\mathcal{L}_{\text{U-PINN}}$ $=\mathcal{L}_{\text{pde}}+\mathcal{L}_{\text{ic}}+\mathcal{L}_{\text{bc}}$ (6)
$=\frac{1}{N_{\text{pde}}}\sum_{i=1}^{N_{\text{pde}}}\left(\frac{\partial\hat{u}_{i}}{\partial t}+\hat{u}_{i}\frac{\partial\hat{u}_{i}}{\partial x}-\nu\frac{\partial^{2}\hat{u}_{i}}{\partial x^{2}}\right)^{2}$
$+\frac{1}{N_{\text{ic}}}\sum_{j=1}^{N_{\text{ic}}}\left(\hat{u}(x_{j},0)+\sin(\pi x_{j})\right)^{2}$
$+\frac{1}{N_{\text{bc}}}\sum_{k=1}^{N_{\text{bc}}}\left(\hat{u}(-1,t_{k})^{2}+\hat{u}(1,t_{k})^{2}\right).$
The semi-supervised S-PINN combines both objectives, with data and physics terms weighted equally:
$\mathcal{L}_{\text{S-PINN}}=\mathcal{L}_{\text{data}}+\mathcal{L}_{\text{pde}}+\mathcal{L}_{\text{ic}}+\mathcal{L}_{\text{bc}}.$ (7)
All models were trained using Adam optimizer followed by L-BFGS optimizer [37] with learning rate $10^{-3}$. Dataset sizes range from $10^{3}$ to $5\times 10^{3}$ training points for PINNs and up to $8\times 10^{3}$ for the NN. Data and code are adapted from [38].
0.A.2 Results
Both PINNs achieve consistently lower test losses at lower carbon footprints compared to the supervised NN (Fig. 1-b). While the NN has the lowest carbon footprint for a fixed number of epochs and data points, achieving a given accuracy level is up to $10$ times more carbon-efficient with the PINNs. The S-PINN, which combines data and physics, dominates the Pareto front, suggesting that the right balance between data and physics yields the most favourable trade-off between predictive performance and carbon cost.
Appendix 0.B Incompressible Shear Flow Dataset
This experiment is based on data from The Well [26], consisting of 2D incompressible Navier–Stokes dynamics with an additional passive tracer:
$\partial_{t}\mathbf{u}+\nabla p-\nu\nabla^{2}\mathbf{u}$ $=-\mathbf{u}\cdot\nabla\mathbf{u},$ (8)
$\partial_{t}s-D\nabla^{2}s$ $=-\mathbf{u}\cdot\nabla s,$ (9)
$\nabla\cdot\mathbf{u}$ $=0.$ (10)
Here $\mathbf{u}=(u_{x},u_{y})$ is the velocity components, $p$ is pressure satisfying $\int p=0$, $s$ is the tracer, $\nu$ is viscosity, and $D$ is diffusivity. The PDE parameters are linked to Reynolds and Schmidt numbers as $\nu=1/\text{Re}$ and $D=\nu/\text{Sc}$.
The system is initialized with a horizontal shear flow:
$u_{x}(y)=\tanh\left(5\frac{y-y_{k}}{n_{\text{shear}}w}\right),$ (11)
where $y_{k}$ is the vertical position of the $k$-th shear layer, $n_{\text{shear}}$ is the number of shear layers, and $w$ is a width parameter that controls how sharply the velocity changes across each layer.
To introduce small perturbations in the system, the vertical velocity field $v_{y}$ contains sinusoidal variations along the $x$-direction, localized at the shear layers. These perturbations are given by:
$u_{y}(x,y)=\sin(n_{\text{blobs}}\pi x)\exp\left(\frac{-25}{w^{2}}|y-y_{k}|^{2}\right).$ (12)
The number of oscillations is controlled by $n_{\text{blobs}}$, while $w$ influences how localized these perturbations are around the shear layers.
The tracer field is initialized to match the shear flow, while the pressure field is initialized to zero everywhere in the domain. Thus, the initial conditions are fully characterized by three parameters: $n_{\text{shear}}\in\{2,4\}$, $n_{\text{blobs}}\in\{2,3,4,5\}$, and $w\in\{0.25,0.5,1.0,2.0,4.0\}$.
As part of preprocessing, we standardize each field using $\hat{x}=(x-\mu)/\sigma$. This normalization balances the contribution of different fields by preventing those with high variance from dominating the loss.
Appendix 0.C Model Configurations
Model configurations are adapted from prior work. U-nets, FNO, and TFNO follow Ohana et al. [26], while UFNO follows Rahman et al. [30]. All models are scaled to 17–20M parameters. Table 3 summarises configurations and parameter counts.
Table 3: Model configurations.
Model Filter size/modes Init dim Blocks/stage Up/down Bottleneck Params ($\times 10^{6}$)
UNet, UNet-CP 33 $48$ 11 44 11 $17.5$
UNet-CN, UNet-CN-CP 77 $42$ 22 44 11 $18.6$
FNO, TFNO $16$ $128$ 44 $19.0$, $19.3$
UFNO $8,12,16$ $128$ 11 33 11 $18.8$
Appendix 0.D Rollout Results
Figs. 8-11 provide rollout prediction examples for each of the four fields. The baseline U-net (UNet and UNet-CP) deteriorate rapidly, accumulating distortions within just a few steps. The modernised variants (UNet-CN, UNet-CN-CP) show marked improvements: UNet-CN captures fine-scale detail but gradually disrupts large-scale dynamics, while UNet-CN-CP better preserves structure at both scales at the cost of high-frequency artifacts. FNO and TFNO maintain dominant large-scale dynamics but exhibit a systematic drift and fail to capture fine-scale variability. UFNO recovers both large- and small-scale structures more effectively, though these gradually smooth out over time.
Figure 8: Example of rollout predictions for horizontal velocity field. Target trajectory (top row) compared with model predictions shown every third step of a 20-step rollout. Models labels are ordered from no bias (top) to strong bias (bottom).
Figure 9: Example of rollout predictions for vertical velocity field. Target trajectory (top row) compared with model predictions shown every third step of a 20-step rollout. Models labels are ordered from no bias (top) to strong bias (bottom).
Figure 10: Example of rollout predictions for pressure field. Target trajectory (top row) compared with model predictions shown every third step of a 20-step rollout. Models labels are ordered from no bias (top) to strong bias (bottom).
Figure 11: Example rollout predictions for the tracer field. Target trajectory (top row) compared with model predictions shown every third step of a 20-step rollout. Models labels are ordered from no bias (top) to strong bias (bottom).